OpenStudy (anonymous):

For the inverse variation equation of xy = 19, complete the statement, y________as x decreases. increases decreases remains constant

OpenStudy (anonymous):

solve for x or y so x=19/y y=19/x plug in values where x decreases to observe y

OpenStudy (anonymous):

for observing y, better off using y=19/x

OpenStudy (anonymous):

As x decreases or becomes smaller and smaller, what happens to y?

OpenStudy (anonymous):

it increases right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks:)